OpenStudy (anonymous):

Find the volume of the sphere to the nearest whole number. Use p = 3.14. r = 4.3 m

OpenStudy (shamil98):

you mean pi = 3.14

OpenStudy (shamil98):

\[v = (4/3) {\pi}r^3\]

jigglypuff314 (jigglypuff314):

volume of sphere = 4/3 pi r^3 so you'll get = 4/3 (3.14) (4.3)^3 =

OpenStudy (studygurl14):

\[V = \frac{ 4 }{ 3 } \pi r^{3}\] \[V = \frac{ 4 }{ 3 } (3.14) (4.3)^{3}\]

OpenStudy (anonymous):

i get 77.4

jigglypuff314 (jigglypuff314):

remember that the (4.3) is cubed :) so (4.3)^3 = (4.3) (4.3) (4.3)

OpenStudy (anonymous):

i meant 332.8

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

thank you for helping me

jigglypuff314 (jigglypuff314):

glad we could help :) Have a Nice Day! And Happy New Year :D